$SPAD/src/input images1.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- Color gallery page 1
)read tknot
-- draw a (15,17) torus knot
torusKnot(15,17, 0.1, 6, 700)

———-

References^{[1] nothing}

4